# ProACT: An Augmented Reality Testbed for Intelligent Prosthetic Arms

### v1

@@ Abstract
Upper-limb amputees face tremendous difficulty in operating dexterous powered prostheses. Previous work has shown that aspects of prosthetic hand, wrist, or elbow control can be improved through “intelligent” control, by combining movement-based or gaze-based intent estimation with low-level robotic autonomy. However, no such solutions exist for whole-arm control. Moreover, hardware platforms for advanced prosthetic control are expensive, and existing simulation platforms are not well-designed for integration with robotics software frameworks. We present the Prosthetic Arm Control Testbed (ProACT), a platform for evaluating intelligent control methods for prosthetic arms in an immersive (Augmented Reality) simulation setting. Using ProACT with non-amputee participants, we compare performance in a Box-and-Blocks Task using a virtual myoelectric prosthetic arm, with and without intent estimation. Our results show that methods using intent estimation improve both user satisfaction and the degree of success in the task. To the best of our knowledge, this constitutes the first study of semi-autonomous control for complex whole-arm prostheses, the first study including sequential task modeling in the context of wearable prosthetic arms, and the first testbed of its kind. Towards the goal of supporting future research in intelligent prosthetics, the system is built upon on existing open-source frameworks for robotics, and is available at https://armlabstanford.github.io/proact.
@@ 1 Introduction
Prosthetic arms are typically controlled via signals acquired through electromyography (EMG) from the residual limb [1, 2]. As the number of input signals is inherently limited, these devices either have limited mechanical functionality, or have more degrees-of-freedom (DoF) than can be simultaneously controlled by their users. Dissatisfaction with functionality is a major reason for rejection and abandonment [3], often within a month of use [4]. Rejection rates vary between $30-80\%$, and are the highest for limb loss near the shoulder [5, 6]. Many individuals with high-level and/or bilateral limb loss wish to use prosthetic arms, but find that, in their current state, these arms introduce more challenges than they alleviate [7, 5].
@@ 1.1 Intelligent prosthetics
In order to provide intuitive high-DoF operation with the low-DoF EMG input available, other sources of information must be utilized to automate low-level functions [8]. In the present work, as in the literature, we use the term “intelligence” as a shorthand for such context-informed shared autonomy, as opposed to direct joint-level control.
@@
HoloPHAM [28] Training MPL EMG Immersive AR Unity not available
Virtual BBT [29] Training Human avatar EMG Immersive VR Unity not available
@@
Unfortunately, the problem of whole-arm control, as in the case of shoulder disarticulation or forequarter amputation, receives little attention in the engineering literature. Krausz et al. (2020) characterized the coordination between gaze and $14$-channel EMG data from non-amputees, showing that gaze could be used with even a subset of EMG channels to predict grasping motions. Using shoulder adduction/abduction and biting motions as input, Togo et al. (2022) demonstrated intelligent control of a custom-built 4-DoF whole-arm prosthesis [30] using gaze tracking to identify the target, and computer vision to plan reach and grasp motions. However, neither of these works was followed by studies extending the proposed methods to the control of higher-DoF wearable arms. Research on intelligent control methods for whole arms can benefit people with all levels of amputation.
@@ 1.2 Enhancing EMG inputs
Targeted Muscle Reinnervation (TMR) [31] surgically increases the number and quality of EMG signals available from the residual limb, bringing myoelectric control within the realm of possibility for whole-arm prosthesis control. However, when successful, it creates at most $2$-$4$ independent channels of simultaneous control inputs [32]; for comparison, the most advanced anthropomorphic prosthesis, the Modular Prosthetic Limb (MPL) [33] has $26$-DoF with $17$ actuators. The feasibility of TMR also depends upon the conditions of limb loss, time elapsed, and surgeon training [34]. Thus, TMR alone cannot make whole-arm prosthetics intuitive to use, yet it is the most prominent development thus far in whole-arm control. Intelligent control can benefit whole-arm prosthetics, both when TMR is possible and when it is not.
@@ 1.3 Immersive AR platforms
Virtual environments are popular in training protocols for prosthesis users [35], as well as in testing new methods of control for both prosthetics [36] and robot teleoperation [37, 38]. Without procuring or fabricating hardware, researchers may study a wide range of devices and conditions in simulation using a single versatile setup, quickly and economically. Immersive platforms provide more realistic results pertaining to human interaction, as compared to visualizations on flat screens [39]. Virtual Reality (VR) systems such as ArmSym [27] and the Virtual Integration Environment or VIE [25] take advantage of this. The HoloPHAM platform [28] uses AR for direct EMG control training. Unlike VR, as mentioned in Section 1.1, AR is useful even as a component of real prosthetic systems, relaying information sensed from the environment (such as object recognition), information sensed from the user (such as gaze), controller decisions, and menu options. An AR evaluation platform for intelligent prostheses would integrate such features with the extensive software infrastructure that exists for robot control. To the best of our knowledge (summarized in Table 1), such a platform does not exist.
@@ 1.5 Problem Statement
The broader research question to be investigated is whether users of powered prosthetic arms – with all levels of amputation – can perform manipulation tasks better and more comfortably, if direct control is augmented with the use of passively revealed cues from the user, sensing and interpretation of the scene and task context, autonomous control of some aspects of joint motion, and AR feedback. In order to address this question, an evaluation platform is needed which: (1) is cost-efficient, time-efficient, and versatile, accommodating studies of partial as well as complete limb loss, with amputee as well as non-amputee participants, (2) is reasonably physically accurate, (3) is reliable for behavioral studies (limiting distortion of participant behavior), (4) is accessible to robotics researchers (compatible with existing task and motion planning frameworks, computer vision libraries, intention prediction models etc. developed for robotics), and (5) is compatible with a variety of input modalities (EMG, joysticks, cameras, gaze trackers) and feedback modalities (tactile actuators, speakers, and a visual display).
@@ 1.6 Contributions
    a user study of intelligent control for a high-DoF wearable prosthetic arm, made possible by the above platform, and
@@
    the inclusion of sequential modeling of a task, as an exploratory component of the study.
Details of the design and implementation of the testbed and the user study are provided in Section 2 below. Results from the study are presented in Section 3 and discussed in Section 4, together with limitations and possible modifications. Section 5 summarizes conclusions from the present work and possible future applications.
@@ 2 Methods
The details of the platform are given in Section 2.1. Section 2.2 explains the design of the evaluation task in light of existing tests. The control methods used in the study are described in Section 2.3. Finally, Section 2.4 reports the design of the user study.
@@ 2.1 System architecture
The platform (Fig. (1a)) is based on ROS Noetic, running on an Ubuntu 20.04 machine (Intel Core i9-12900K CPU, $64$ GB RAM). A real-time dynamics simulation ($1$ kHz update rate, $1$ ms time step) in Gazebo [^1] Unity was not used for dynamics, since it prioritizes visual realism and stability over physical accuracy (By default, Unity and Gazebo rely on the physics engines PhysX and ODE or Open Dynamics Engine respectively [43]). is displayed to participants (Fig. (1b)) at $60$ Hz, via a mixed-reality application written in Unity 2020.3 and deployed to a Microsoft HoloLens 2 Head-Mounted Display (HMD).
Communication: The open-source ROS# libraries [44] are used to communicate between ROS and Unity. An OyMotion EMG armband worn on the forearm sends control inputs over a Bluetooth Low Energy (BLE) connection. This allows for quick setup (no individual electrode placement), while approximating the few sequential control inputs available to a shoulder disarticulation amputee. We use the recently-released LibEMG libraries [45] to classify $5$ discrete hand gestures [46] via Linear Discriminant Analysis (LDA).
Localization: The base of the virtual prosthetic arm tracks motion capture markers attached to a shoulder brace (Fig. (1c)). At the start of each trial, the transform between the Unity and motion capture reference (world) frames is found using the HMD, which is tracked in both systems (via built-in localization and attached markers, respectively). The box and blocks are displayed at a fixed height measured from the shoulder in an upright posture, at a fixed horizontal displacement from the reference board. Having thus located the Gazebo world frame in the physical space, all virtual objects are displayed in their correct poses relative to the Unity world.
Device: The default arm used in the platform is the Modular Prosthetic Limb (MPL), a research platform for sophisticated anthropomorphic whole-arm prostheses. The modern version of the arm, unlike the online model [47], includes shoulder flexion [33], hence we add a DoF at the base. The online $26$-DoF model, unlike the $26$-DoF model described in the chapter [33], has a $20$-DoF hand (rather than $19$). The addition of the shoulder flexion joint results in a $7$-DoF arm and $20$-DoF hand. Such a high-DoF prothesis, impossible to control completely with $2$-$4$ sequential inputs, is more suitable for studying the impact of intent estimation than one which may have a one-to-one mapping of input channels, but may not have sufficiently high DoF to perform complex tasks.
@@ 2.2 Task design
The Box-and-Blocks Test [48] is a well-established measure of upper-limb dexterity for stroke patients in recovery. It has been modified for specific conditions such as Parkinson’s disease [49], cerebral palsy [50, 51], fibromyalgia [52], and prosthesis use [53, 54, 55]. The task requires participants to transport blocks from one side of a box to another, over a partition. The partition and walls of the box ensure that some extent of complex manipulation is demanded. Virtual versions of the test have been developed for many of these conditions as well [56, 57, 29]. In the original test, $150$ blocks are heaped on one side and the score is defined by the number of blocks moved in a fixed time [58]. Modifications for the evaluation of prosthesis use in particular, the mBBT [54] and tBBT [55], decrease the number of blocks, and specify the initial positions, place locations, and sequence of interaction. The score, in these tests, is defined by the time taken to move all the blocks. The controlled sequence reduces variability between subjects introduced by choice, and controlled place locations make it possible to study the entire task including precision pick and place.
@@
Modifications to task: We make the following further modifications: (1) The size of blocks is increased to mitigate visual occlusion by the hand. (2) The number of blocks is further decreased and spacing increased, so the box dimensions remain reachable, with the walls continuing to constrain reaching. (3) As real-time, realistic contact simulation is difficult, cylindrical blocks are used, where grasp shape is agnostic to approach direction. (4) Numbered blocks of two different colors are used. Color and number are used as a proxy for the semantics of a sequential task, with only the color being taken into account as described in Section 2.3. Similarly to the mBBT and tBBT, participants are instructed to move the blocks in a specific order, but the initial configurations are counterbalanced over all possible arrangements. This is necessary because the difficulty of reaching different regions of the box varies widely and it is infeasible to get data for all blocks in all trials (as explained below).
@@
Modifications to format: Various versions of the test require participants to sit [58] or stand [55], and score performance based on either the number of blocks moved in a fixed period of time [58] or the time taken to move all the blocks [55]. For the present context, these parameters were determined based on the following observations from pilot studies: First, when standing and moving freely, pilot participants were able to achieve nearly identical performance in all methods by using whole-body motion for reaching. Seating participants forces them to operate the MPL arm, making comparisons possible. Second, mistakes reflect the quality of control, but affect subsequent actions. Multiple trials are needed so that incidental mistakes do not result in the loss of all data for some participant-method combinations. Finally, the onset of fatigue has a strong effect on performance and time taken. Extended HMD use also leads to slight headaches, eye strain, heating of the device, and battery depletion. The Hololens 2 battery lasts for two hours of continuous operation. Participants’ skill in the task varies widely, some spending $5$-$10$ minutes per block with the more difficult methods. Accounting for the demonstration and training phase, and the need for multiple trials with all methods within a single session of reasonable duration, the duration of each trial is limited to $5$ minutes.
@@
Anthropometric adjustment: The MPL is a large device. Scaling to participant height would require editing of not only geometric but inertial parameters, which is not realistic (real prosthetic arms have weight and dimensions determined by functionality, design and industrial standardization). Hence, rather than scale the arm or the BBT, the table height is set at $0.75$ m below the participant’s shoulder before starting each trial. This distance was determined during pilots, such that it would be possible to comfortably reach all the blocks.
@@ 2.3 Feedback and control
Two hand poses are pre-programmed as “open” and “closed”, the closed pose being a precision circular grasp approaching the block diameter. The Gazebo grasp plugin [59] is used to rigidly attach objects to the fingertips based on proximity. Participants are made aware of this implementation. Distinct sounds are played through the HMD when contact is made or lost. The $5$ hand gestures are mapped to the movements of the $7$ joints of the MPL arm, and the opening and closing of the MPL hand. These movements are referred to below as HO (hand open), HC (hand close), WF (wrist flex), WE (wrist extend) and NM (no motion). Four methods of control are compared (Fig. 2), spanning the spectrum between direct and assisted control.
@@
Gaze-assisted control (C): The participant’s gaze (tracked by the HMD) is used to infer a target, as shown in Fig. (3), which is highlighted in AR. Participants use WF to “lock” the selection and WE to execute motion towards the selected object. For placing, a dark marker tracking gaze appears on the target side of the box. WF locks the position of the spot and WE executes motion towards it. Motion is based on trajectories generated by sample-based planners selected by the Open Motion Planning Library (OMPL), with asynchronous execution, so that NM immediately stops the motion. A new motion plan is attempted every time a change to WE is detected. If WE is held, the existing plan is used. Participants are made aware that, if the base of the arm is moved during execution, a new planning action (NM-WE) might be required. The selection marker changes color upon locking and unlocking the gaze-based selection. When the goal is unreachable, the planner sets an intermediate goal.

### v2

@@ Abstract
Upper-limb amputees face tremendous difficulty in operating dexterous powered prostheses. Previous work has shown that aspects of prosthetic hand, wrist, or elbow control can be improved through “intelligent” control, by combining movement-based or gaze-based intent estimation with low-level robotic autonomy. However, no such solutions exist for whole-arm control. Moreover, hardware platforms for advanced prosthetic control are expensive, and existing simulation platforms are not well-designed for integration with robotics software frameworks. We present the Prosthetic Arm Control Testbed (ProACT), a platform for evaluating intelligent control methods for prosthetic arms in an immersive (Augmented Reality) simulation setting. We demonstrate the use of ProACT through preliminary studies, with non-amputee participants performing an adapted Box-and-Blocks task with and without intent estimation. We further discuss how our observations may inform the design of prosthesis control methods, as well as the design of future studies using the platform. To the best of our knowledge, this constitutes the first study of semi-autonomous control for complex whole-arm prostheses, the first study including sequential task modeling in the context of wearable prosthetic arms, and the first testbed of its kind. Towards the goal of supporting future research in intelligent prosthetics, the system is built upon on existing open-source frameworks for robotics, and is available at https://arm.stanford.edu/proact.
@@ 1 Introduction
Prosthetic arms are typically controlled via signals acquired through electromyography (EMG) from the residual limb [1, 2]. As the number of input signals is inherently limited, these devices either have limited mechanical functionality, or have more degrees-of-freedom (DoF) than can be simultaneously placed under direct independent control. Dissatisfaction with functionality is a major reason for rejection and abandonment [3], often within a month of use [4]. Rejection rates vary between $30-80\%$, and are the highest for limb loss near the shoulder [5, 6]. Many individuals with high-level and/or bilateral limb loss wish to use prosthetic arms, but find that, in their current state, these arms introduce more challenges than they alleviate [7, 5].
@@ 1.1 Intelligent prosthetics
In order to provide intuitive high-DoF operation with the low-DoF EMG input available, other sources of information may be utilized to automate low-level functions [8]. In the present work, as in the literature, we use the term “intelligence” as a shorthand for such context-informed shared autonomy, as opposed to direct joint-level control.
@@
HoloPHAM [28] User training MPL EMG Immersive AR Unity not available
Virtual BBT [29] User training Human avatar EMG Immersive VR Unity not available
@@
However, it is in whole-arm control, as in the case of shoulder disarticulation or forequarter amputation, that users are most often dissatisfied with conventional myoelectric control methods [5, 6]. Unfortunately, this problem receives little attention in the engineering literature. Krausz et al. (2020) characterized the coordination between gaze and $14$-channel EMG data from non-amputees, showing that gaze could be used with even a subset of EMG channels to predict grasping motions. Using shoulder adduction/abduction and biting motions as input, Togo et al. (2022) demonstrated intelligent control of a custom-built 4-DoF whole-arm prosthesis [30] using gaze tracking to identify the target, and computer vision to plan reach and grasp motions. However, neither of these works was followed by studies extending the proposed methods to the control of higher-DoF wearable arms.
As the needs, capabilities, and preferences of myoelectric prosthesis users vary widely [31], we hypothesize that studying intelligent control methods for whole arms, where direct control inputs are the most limited, has the potential to give rise to insights and techniques benefiting people with all levels of amputation in the future111along similar lines to the widespread adoption of predictive text, originally developed for use with disabilities [32]. As with text prediction, the expected variation in user preferences for intelligent prostheses (even within each level of amputation) underscores the need for large and detailed user studies..
@@ 1.2 Enhancing EMG inputs
Targeted Muscle Reinnervation (TMR) [33] surgically increases the number and quality of EMG signals available from the residual limb, bringing myoelectric control within the realm of possibility for whole-arm prosthesis control. When successful, it creates upto $2$-$4$ independent channels of simultaneous control inputs [34], which can be used with advanced pattern recognition techniques to perform coordinated joint motions [34]. The feasibility and outcome of TMR depend upon the conditions of limb loss, time elapsed, and surgeon training [35]. Meanwhile, the most advanced anthropomorphic prosthesis, the Modular Prosthetic Limb (MPL) [36] has $26$-DoF with $17$ actuators. Thus, TMR alone (in spite of being the most prominent development thus far in whole-arm control) cannot make whole-arm prosthetics intuitive to use in all cases.
@@ 1.3 Immersive AR platforms
Virtual environments are popular in training protocols for prosthesis users [37], as well as in testing new methods of control for both prosthetics [38] and robot teleoperation [39, 40]. Without procuring or fabricating hardware, researchers may study a wide range of devices and conditions in simulation using a single versatile setup, quickly and economically.
Immersive platforms provide more realistic results pertaining to human interaction, as compared to visualizations on flat screens [41]. Virtual Reality (VR) systems such as ArmSym [27] and the Virtual Integration Environment or VIE [25] take advantage of this. The HoloPHAM platform [28] uses AR for direct EMG control training. Unlike VR, as mentioned in Section 1.1, AR is useful even as a component of real prosthetic systems, relaying information sensed from the environment (such as object recognition), information sensed from the user (such as gaze and contralateral hand motion), controller decisions, and menu options. AR devices are also less likely to induce simulator sickness during long studies [42, 43], and in ambulatory tasks, give participants the confidence to move around the room naturally. Given the extensive software infrastructure that exists for the control of autonomous robotic manipulators, integrating it with the benefits of AR into a single platform would enable behavioral studies with intelligent prostheses across a broad scope. To the best of our knowledge (summarized in Table 1), such a platform does not exist.
@@ 1.5 Problem Statement
The broader research question to be investigated is whether users of powered prosthetic arms can perform manipulation tasks better and more comfortably, if direct control is augmented with the use of passively revealed cues from the user, sensing and interpretation of the scene and task context, autonomous control of some aspects of joint motion, and AR feedback. In order to address this question, an evaluation platform is needed which: (1) is cost-efficient, time-efficient, and versatile, accommodating studies of partial as well as complete limb loss, with amputee as well as non-amputee participants, (2) is reasonably physically accurate, (3) is informative for behavioral studies (limiting distortion of participant behavior), (4) is accessible to robotics researchers (compatible with existing task and motion planning frameworks, computer vision libraries, intention prediction models and other algorithms developed for robotics), and (5) is compatible with a variety of input modalities (EMG, joysticks, cameras, gaze trackers) and feedback modalities (tactile actuators, speakers, and a visual display).
@@ 1.6 Contributions
    two exploratory user studies of intelligent control for a high-DoF wearable prosthetic arm, demonstrating the usability of the above platform for non-amputees, and
@@
    preliminary data suggesting possible advantages of intelligent control design for whole-arm prostheses, and highlighting concerns relevant to the design of such interfaces.
Details of the design and implementation of the testbed and the user studies are provided in Section 2. Results are presented in Section 3 and discussed in Section 4, together with limitations and possible modifications. Section 5 summarizes outcomes from the present work and future directions.
@@ 2 Methods
The details of the platform are given in Section 2.1. Section 2.2 explains the design of the evaluation task in light of existing tests. The control methods used in the study are described in Section 2.3. Finally, Section 2.4 reports the design of the user studies.
@@ 2.1 System architecture
The platform (Fig. (1a)) is based on ROS Noetic, running on an Ubuntu 20.04 computer (Intel Core i9-12900K CPU, $64$ GB RAM). A real-time dynamics simulation ($1$ kHz update rate, $1$ ms time step) in Gazebo [^2] Unity was not used for dynamics, since it prioritizes visual realism and stability over physical accuracy (By default, Unity and Gazebo rely on the physics engines PhysX and ODE or Open Dynamics Engine respectively [47]). is displayed to participants (Fig. (1b)) at $60$ Hz, via a mixed-reality application written in Unity 2020.3 and deployed to a Microsoft HoloLens 2 Head-Mounted Display (HMD). For localizing the arm in space, a Windows 10 computer (Intel Core i9-9900KS CPU, $64$ GB RAM) transmits motion capture data from an $8$-camera OptiTrack system over a 5G WiFi network. The open-source ROS# libraries [48] are used to communicate between ROS and Unity.
Control: An OyMotion EMG armband worn on the forearm sends control inputs over a Bluetooth Low Energy (BLE) connection. This allows for quick setup (no individual electrode placement), while approximating the few sequential control inputs that might be available to a shoulder disarticulation amputee. We use the LibEMG libraries [49] to classify $5$ discrete hand gestures via Singular-Value Decomposition-based Linear Discriminant Analysis (LDA), publishing classes as ROS messages [^3] A message is published only if the classifier returns the same output three times consecutively., subscribed to with a queue length of $1$ (to use the latest input). A Mean absolute value, zero crossings, slope sign change and waveform length are extracted from data acquired at $1000$ Hz, using a sliding window size of $100$ samples with a $50$-sample increment between windows [50].
Localization: The base of the virtual prosthetic arm tracks motion capture markers attached to a shoulder brace (Fig. (1c)), with the ROS tf2 transform library being used to manage coordinate frame transformations over time. At the start of each trial, the transform between the Unity and motion capture reference (world) frames is found using the HMD. The HMD is tracked in both systems (via built-in localization and attached markers, respectively). The box and blocks are displayed at a fixed height measured from the shoulder in an upright posture, at a fixed horizontal displacement from the reference board. Having thus located the Gazebo world frame in the physical space, all virtual objects are displayed in their correct poses relative to the Unity world (resetting takes under a minute).
Device: The default arm used is the Modular Prosthetic Limb (MPL), a research platform for sophisticated anthropomorphic whole-arm prostheses. The modern version of the arm, unlike the online model [51], includes shoulder flexion [36], hence we add a DoF at the base. The online $26$-DoF model, unlike the $26$-DoF model described in the chapter [36], has a $20$-DoF hand (rather than $19$). The addition of the shoulder flexion joint results in a $7$-DoF arm and $20$-DoF hand. Such a high-DoF prothesis, impossible to control completely with $2$-$4$ sequential inputs, is more suitable for studying the impact of intent estimation than one which may have a one-to-one mapping of input channels, but may not have sufficiently high DoF to perform complex tasks.
@@ 2.2 Task design
The Box-and-Blocks Test [52] is a well-established measure of upper-limb dexterity for stroke patients in recovery. It has been modified for specific conditions such as Parkinson’s disease [53], cerebral palsy [54, 55], fibromyalgia [56], and prosthesis use [57, 58, 59]. The task requires participants to transport blocks from one side of a box to another, over a partition. The partition and walls of the box ensure that some extent of complex manipulation is demanded. Virtual versions of the test have been developed for many of these conditions [60, 61, 29]. In the original test, $150$ blocks are heaped on one side and the score is defined by the number of blocks moved in a fixed time [62]. Modifications to evalute prosthesis use in particular, the mBBT [58] and tBBT [59], decrease the number of blocks, and specify the initial positions, place locations, and sequence of interaction. The score is defined by the time taken to move all the blocks. The controlled sequence reduces variability between subjects introduced by choice, and controlled place locations make it possible to study the entire task including precise actions.
@@
Modifications to task: We make the following further modifications: (1) The size of blocks is increased to mitigate visual occlusion by the hand. (2) The number of blocks is further decreased and spacing increased, so the box dimensions remain reachable, with the walls continuing to constrain reaching. (3) Cylindrical blocks are used, where grasp shape is agnostic to approach direction. (4) Numbered blocks of two different colors are used. Color and number are used as a proxy for the semantics of a sequential task, with only the color being taken into account as described in Section 2.3. Similarly to the mBBT and tBBT, the order is specified, but the initial configurations are counterbalanced over their possible arrangements, as the difficulty of reaching different regions of the box varies widely and it is infeasible to get data for all blocks in all trials (as explained below).
@@
Modifications to format: Various versions of the test require participants to sit [62] or stand [59], and score performance based on either the number of blocks moved in a fixed period of time [62] or the time taken to move all the blocks [59]. For the present context, these parameters were determined based on the following observations from pilot studies: First, when standing and moving freely, participants were able to achieve nearly identical performance in all methods by using whole-body motion for reaching. Seating them forces them to operate the MPL arm, making comparisons possible. Second, mistakes reflect the quality of control, but affect subsequent actions. Multiple trials are needed so that incidental mistakes do not impact all data for some participant-method combinations. Finally, the onset of fatigue has a strong effect on performance. Extended HMD use also leads to slight headaches, eye strain, heating of the device, and battery depletion. The Hololens 2 battery lasts for $2$-$3$ hours of continuous operation. Participants’ skill in the task varies widely, some spending $5$-$10$ minutes per block with the more difficult methods. Accounting for the demonstration and training phase, and the need for multiple trials within a single session of reasonable duration, each trial is limited to $5$ minutes.
@@
Anthropometric adjustment: The MPL is a large device. Scaling to participant height would require editing of not only geometric but inertial parameters, which is not realistic (real prosthetic arms have weight and dimensions determined by functionality, design and industrial standardization). Hence, rather than scale the arm or the BBT, the table height is set at $0.43$ m below the measured location of the MPL shoulder before starting each trial. This distance was determined during pilots, such that it would be possible to comfortably reach all the blocks. The partition is aligned with the participant’s sagittal plane using the reference board, which can be slightly corrected based on the participant’s reporting if necessary. The horizontal distance between the participant and the box was adjusted by participants in Study 1 (to be as close as possible without intersecting their legs), and was fixed at $0.6$ m to the box center in Study 2 (for better consistency across participants, considering the fixed size of the arm).
@@ 2.3 Feedback and control
Two hand poses are pre-programmed as “open” and “closed”, the closed pose being a precision circular grasp approaching the block diameter. In place of simulating contact physics, the Gazebo grasp plugin [63] is used to rigidly attach objects to the fingertips based on proximity. Participants are made aware of this implementation. Distinct sounds are played through the HMD when contact is made or lost. The $5$ hand gestures are mapped to the movements of the $7$ joints of the MPL arm, and the opening and closing of the MPL hand. These movements are referred to below as HO (hand open), HC (hand close), WF (wrist flex), WE (wrist extend) and NM (no motion). Four methods of control are compared (Fig. 2), spanning the spectrum between direct and assisted control.
@@
Gaze-assisted control (C): The participant’s gaze (tracked by the HMD , within a $1.5\degree$ error) is used to infer a target, as shown in Fig. (3), which is highlighted in AR. Participants use WF to “lock” the selection and WE to execute motion towards the selected object. For placing, a dark marker tracking gaze appears on the target side of the box. WF locks the position of the spot and WE executes motion towards it. Motion is based on trajectories generated by sample-based planners selected by the Open Motion Planning Library (OMPL), with asynchronous execution, so that NM immediately stops the motion. A new motion plan is attempted every time a change to WE is detected. If WE is held, the existing plan is used. Participants are made aware that, if the base of the arm is moved during execution, a new planning action (NM-WE) might be required. The selection marker changes color upon locking and unlocking the gaze-based selection. When the goal is unreachable, the planner sets an intermediate goal.